# Learning Contrastive Feature Representations for Facial Action Unit Detection

### v6

@@ 1 Introduction
Facial Action Coding System (FACS) [1] is a widely employed approach for facial expression coding, which introduced a set of 44 facial action units (AUs) that establish a link between facial muscle movements and facial expressions [2]. AU detection is a multi-label binary classification problem. Each AU represents a kind of label, and activated and inactivated AUs are respectively labeled 1 and 0. Nowadays, AU detection [3, 4] has emerged as a computer vision technique to automatically identify AUs in videos or images. This advancement holds promise for diverse applications, including human-computer interaction, emotion analysis, car driving monitoring et al.
In recent years, supervised frameworks for AU detection have predominantly relied on binary cross-entropy loss to align predicted and ground truth labels [5, 6, 7]. While this approach is common, it often places unnecessary emphasis on strict label alignment, increasing the risk of overfitting by demanding excessive model capacity. Additionally, these methods tend to focus on encoding pixel-level facial features, overlooking the fact that capturing subtle differences in different AUs’ activation is sufficient. As a result, such approaches unnecessarily increase model complexity, limiting the generalization ability of the model.
@@
As a result, representation learning approaches based on discriminative contrastive learning have emerged and shown significant potential [8] for classification tasks. As shown in Fig. 1, this approach aims to maximize the consistency of positive samples that are semantically similar while minimizing the consistency of negative samples that are semantically dissimilar. By optimizing the contrastive loss, the mutual information between variables is either explicitly or implicitly optimized [9]. However, the direct integration of contrastive learning into AU detection encounters two principal challenges. The first challenge is the class imbalance issue of each AU type [10]. Most AU classes manifest a substantial imbalance between majority and minority class samples, with inactivated AUs vastly outnumbering their activated counterparts. This imbalance issue influences the classification decision threshold, potentially resulting in the direct classification of the majority of AUs as the negative class. The second challenge pertains to the noisy and false label problem [11]. Due to the subtle feature differences when AUs activate leading to the inherent difficulty in expertly labeling AUs, existing AU datasets may exhibit a few noisy and false labels, significantly undermining the overall fitting ability of models.
To address these challenges, we propose a discriminative contrastive learning framework inspired by self-supervised learning techniques for learning highly semantic image representations [8]. This framework encodes differential information between AUs from each sample pair, offering an alternative to the traditional pixel-level feature learning methods commonly employed in AU detection. The resulting loss, termed AUNCE, is conceptually similar to the widely used InfoNCE loss. Moreover, to tackle the class imbalance issue in AU detection, we adopt a negative sample re-weighting strategy inspired by [7]. This strategy addresses class imbalance by assigning varying levels of importance to individual samples during training, achieved by adjusting the weights of their loss contributions. Additionally, to mitigate the effects of noisy and false AU labels, inspired by the sample-mixing concept proposed in [12], we introduce a positive sample sampling strategy. This strategy enhances supervision signals by incorporating self-supervised signals, achieved by defining sampling ratios for different types of positive sample pairs. This ensures that more robust positive samples are derived through data augmentation combined with self-supervised signal. The primary contributions of our work are summarized as follows:
@@
    We introduce a positive sample samplying strategy to mitigate the effects of noisy and false AU labels.
@@ 2.1 Discriminative contrastive learning
In recent years, discriminative contrastive learning, a subset of deep self-supervised learning methods [13, 14], has witnessed substantial progress. This approach follows a learn-to-compare paradigm, distinguishing between similar (positive) and dissimilar (negative) data points in a feature space. For instance, Oord et al.[9] introduced the InfoNCE loss to optimize feature extraction by maximizing consistency in sequence data, establishing a foundation for contrastive methods. Chen et al.[13] proposed SimCLR, which uses a base encoder and projection head to maximize agreement through InfoNCE. Similarly, MoCo [14] leverages a momentum network to maintain a queue of negative samples for efficient learning. Chen and He [15] addressed collapsing solutions by avoiding negative pairs via a stop-gradient operation. While variations in representation encoders and similarity measures exist, the core principle remains: training a representation encoder $f$ by contrasting similar pairs ($x,x^{+}$) against dissimilar pairs ($x,x^{-}$) to optimize the contrastive loss.
Apart from that, deep contrastive learning not only facilitates the automated extraction of refined feature representations through the comparison of positive and negative sample pairs but also incorporates diverse contrastive losses for optimization. Khosla et al. [16] extended the contrastive learning concept to supervised settings with a supervised contrastive learning loss, aiming to enhance feature expression using labeled data. Additionally, Chen et al. [17] proposed a semi-supervised contrastive learning algorithm, initially pretraining all data through contrastive learning and subsequently transferring knowledge to new models via distillation learning with the aid of labeled data.
@@ 2.2 Contrastive learning for AU detection
Contrastive learning was firstly designed as a pretext task for unsupervised or self-supervised learning and was introduced to the domain of AU detection with limited annotations [18, 19, 20]. For example, Niu et al. [18] introduced a semi-supervised co-training approach that utilized contrastive loss to enforce conditional independence between facial representations generated by two Deep Neural Networks (DNNs). Following this, Zhang et al. [19] introduced a weakly-supervised contrastive learning approach that effectively utilizes coarse-grained activity information. Additionally, Liu et al. [20] introduced a semi-supervised contrastive learning strategy that concurrently combined region learning, contrastive learning, and pseudo-labeling.
@@
However, previous methods primarily employed a combination of weighted cross-entropy loss and triplet loss to constrain the training process, encoding pixel-level information of the entire face associated with AUs. This results in significant redundant information that may dilute the discriminative signals necessary for accurate AU detection. To address this, Sun et al.[23] proposed EmoCo, an unsupervised discriminative contrastive learning framework that uses emotion-based image pairs for AU detection. While EmoCo improves AU detection, it faces challenges due to conflicting AU labels in emotion-based pairs, leading to false positives and false negatives, and its reliance on emotion annotations limits its applicability to datasets with only AU labels. Similarly, Li et al.[11] introduced a self-supervised discriminative contrastive learning approach using the MoCo framework [14] for dynamic AU detection in video datasets. However, this method is limited to dynamic AU detection and cannot be directly applied to static image datasets without temporal information.
@@
(a) The first stage of our framework: pretraining
(b) The second stage of our framework: linear evaluation
Figure 2: Overview of the training pipeline of our discriminative contrastive learning framework. The training process consists of two stages: pretraining and linear evaluation. In the first stage, the feature encoder is pretrained by the AUNCE loss. In the second stage, the learned feature representations are utilized for AU detection, with a linear classifier trained using the weighted cross-entropy loss (WCE). The linear classifier serves to evaluate performance without heavily influencing feature learning.
@@ 3.1 Objectives
To address these challenges, Section 3 provides detailed solutions. In Section 3.2, we propose a discriminative contrastive learning framework that directs the encoder to capture differential features between AUs from each sample pair, thus moving away from conventional pixel-level feature learning. To combat the class imbalance issue, Section 3.3 introduces a negative sample re-weighting strategy, drawing inspiration from, which prioritizes underrepresented AUs and facilitates the detection of challenging samples. Besides, Section 3.4 presents a positive sample sampling strategy that incorporates augmented supervision signals, which helps the model effectively learn from noisy and false data. Finally, Section 3.5 will introduce the encoder structure we used in our framework.
@@ 3.2 Discriminative contrastive learning framework
where $y_{i}$ denotes the ground truth value of occurrence (i.e., 0 or 1) of the $i$-th AU, and $\hat{y}_{i}$ is the corresponding predicted occurrence probability. This loss effectively guides the training process, encouraging the model to allocate higher probabilities to the correct class labels and lower probabilities to incorrect ones. $w_{i}$=$\frac{(1/r_{i})n_{au}}{\sum_{i=1}^{n_{au}}(1/r_{i})}$ is the weight of the $i$-th AU determined by its occurrence frequency for balanced training. It permits reducing the impact of loss values associated with those AUs that are more commonly activated in the training set. $r_{i}$ represents the occurrence frequency of the $i$-th AU, and $n_{au}$ is the total number of AUs.
@@
To address this challenge, we introduce a discriminative contrastive learning framework for AU detection. As shown in Fig. 2, the training process of our framework consists of two stages: pretraining and linear evaluation. During the pretraining stage shown in Fig. 2(a), our framework includes not only the feature encoder but also a positive sample mining module and a negative sample re-weighting module, which together assist the training process of the encoder. The former integrates self-supervised signals with supervised signals to address the issue of noisy labels arising from subtle differences in AUs. The latter is introduced to tackle challenges related to class imbalance and hard negative sample mining. Detailed discussions of these two modules are provided in Sections 3.3 and 3.4, respectively.
@@
The adoption of ${L}_{aunce}$ over ${L}_{wce}$ during the pretraining phase is motivated by several key considerations: 1. The weighted cross-entropy loss is computed based on the sample and its corresponding label $y$, requiring the encoding of pixel-level information in order to align with the label. In contrast, the contrastive loss does not involve explicit labels but is computed based on the feature difference values between each pair of samples. This incentivizes the encoder to capture the differential information between AUs from each sample pair, without the need to focus on pixel-level information extraction of individual samples, reducing the borden of encoder. 2. The weighted cross-entropy loss imposes constraints on the absolute values of the predicted probabilities. However, for classification tasks, this is often unnecessary and can lead to overfitting as well as demanding higher fitting capability from the encoder. The contrastive loss focuses on constraining the similarity difference between positive and negative sample pairs without explicitly specifying their exact values. This is because the minimal value of $\mathcal{L}_{aunce}$ is 0, which is achieved when $\text{sim}((f(x),f(x^{+}))-\text{sim}(f(x),f(x^{-}_{j}))\rightarrow+\infty$ for all negative sample index $j\in{1,2,\cdots,N}$.
In evaluating the representations obtained from the pretraining process, we adhere to the established linear evaluation protocol, as commonly employed in previous studies [9, 13]. As illustrated in Fig. 2(b), a linear classifier is trained atop the encoder with fixed parameters, utilizing ${L}_{wce}$ during the training phase. It is important to note, however, that this linear classifier consists of a simple fully connected layer and serves primarily for performance evaluation rather than significantly impacting the feature learning process. Therefore, the model’s core capability is largely determined by the effectiveness of AUNCE in guiding the feature extraction.
3.3 Negative sample re-weighting strategy
@@
The key difference between our weighting method and HCL [25] lies in the absence of mean normalization in Eq. (3), allowing for more effective handling of class imbalance in AU detection. Specifically, when the anchor sample $x$ belongs to the majority class, its positive samples (majority class) receive smaller gradient magnitudes to limit parameter updates, while its negative samples (minority class) receive larger gradient magnitudes to emphasize their impact. To achieve this, we set $\beta\geq 1$, increasing the influence of updates caused by minority class samples. Conversely, when $x$ belongs to the minority class, we set $\beta\leq 1$, reducing the influence of majority class samples. By tuning $\beta$, we control the relative importance of positive and negative samples, ensuring the model focuses appropriately on minority class samples and balances gradient magnitudes. Including the mean as a normalization factor would undermine this control, as changes in $\beta$ would simultaneously affect the mean, leading to unpredictable weight adjustments. Thus, the omission of mean normalization in Eq. (3) is a deliberate choice to address class imbalance, differentiating our approach from HCL.
3.4 Positive sample samplying strategy
For a given anchor point, samples sharing the same label with the anchor are referred to as positive examples. Among them, positive examples whose ground truth labels do not match the anchor’s label are termed as false positives, primarily stemming from label noise in the samples. Regarding the third challenge of AU labels tainted by noise and errors, we employ a combined approach that integrates label noisy learning and self-supervised learning techniques to obtain enhanced supervised signals for positive samples. As shown in Fig. 3, this approach mitigates the adverse effects of noisy labels by simultaneously introducing a controlled level of self-supervised signal and supervised signal, thereby enhancing the robustness of model to variations and errors present in the labeled dataset.
@@ 3.4 Positive sample samplying strategy
3.4.1 Supervised Signal
One common technique to address the issue of false data is label smoothing, where a small amount of uncertainty is incorporated into the training process by replacing one-hot labels with smoothed probability distributions. However, it is challenging to directly apply label smoothing to contrastive learning as the computation of contrastive loss does not involve explicit labels. To overcome this challenge, we turn to the memory effect of neural networks. Previous research has shown that neural networks tend to first fit clean samples and then memorize the noise in incorrectly labeled samples. This memory effect is independent of the choice of neural network architecture. The positive samples that are easier fit are those that are closer to the anchor point and have higher similarity. Inspired by this observation, for a given sample $x$ and M positive samples ${x}_{i=1}^{M}$ with noisy labels belonging to the same class, we calculate the similarity scores between the representation of sample $x$ and the representations of all M positive samples. We then select the positive feature representation with the highest similarity to sample $x$:
@@ 3.4.1 Supervised Signal
3.4.2 Self-supervised Signal
In addition to label noisy learning, we incorporate self-supervised learning techniques into the training process. By designing these auxiliary tasks carefully, we can encourage the model to learn more robust and invariant representations, which can help mitigate the effects of label noise and improve generalization performance. Specifically, our second method to obtain positive feature involves performing semantically invariant image augmentation on the anchor point $x$:
@@ 3.4.2 Self-supervised Signal
The primary distinction from data augmentation approaches exemplified by SimCLR lies in the nature of positive examples. In SimCLR, positive examples are augmented versions of the anchor point, implying that two views of the same sample are defined as positives, with each sample representing a unique class, leading to instance-level contrast. Under such circumstances, positive examples do not fully encapsulate the general class information. In contrast, the proposed method defines positive examples as the mean of the same-class samples that are identical to the anchor point, essentially representing the most representative features fo positive samples, which can be interpreted as class prototypes. Compared to instance-level contrast, the feature representations of class prototypes offer more robust information for comparison, thereby better capturing class information. The positive feature representation $f(x_{m}^{+})$ obtained from Eq. (10) represents the class centroid of M positive examples $\{x\}_{i=1}^{M}$. In other words, it synthesizes a virtual representative positive sample.
The aforementioned three methods of obtaining positive examples combine the use of label noisy learning for supervised signals and data augmentation for self-supervised signals. These three types of positive examples are randomly mixed with probabilities $p_{1}$, $p_{2}$, and $p_{3}$ respectively, where $p_{1}+p_{2}+p_{3}=1$. Our approach aims to improve the quality of supervised signals utilized in the training process. By integrating these techniques, we effectively mitigate the impact of noisy and false labels, increase the resilience of models to errors and variations in labeled datasets, and ultimately enhance the generalization ability of models to unseen data.
@@
Input: Mini batch data $\mathcal{R}$, feature representation function $f(\cdot)$, $p_{1}$, $p_{2}$, and $p_{3}$, $\beta$

### v7

@@ 1 Introduction
Facial Action Coding System (FACS) [1] is a widely employed approach for facial expression coding, which introduced a set of 44 facial action units (AUs) that establish a link between facial muscle movements and facial expressions [2]. AU detection is a multi-label binary classification problem. Each AU represents a kind of label, and activated and inactivated AUs are respectively labeled 1 and 0. Nowadays, AU detection [3] has emerged as a CV technique to automatically identify AUs in videos or images. This advancement holds promise for diverse applications, including human-computer interaction, emotion analysis, car driving monitoring et al..
In recent years, supervised learning has dominated AU detection research [4, 5], typically employing binary cross-entropy (BCE) loss for optimization. However, the emphasis of BCE on precise label alignment and pixel-level encoding often leads to overfitting and excessive model complexity, limiting generalization to unseen data. Shifting from pixel-level features to inter-AU differential encoding allows the model to capture AU-specific spatial dependencies by emphasizing relative variations among AUs rather than absolute intensities. This abstraction is theoretically consistent with the Facial Action Coding System (FACS), where facial expressions emerge from the coordinated activation of multiple AUs.
@@
Representation learning based on discriminative contrastive learning has shown substantial potential [6] for classification tasks. As illustrated in Fig. 1, it maximizes similarity among semantically positive samples while minimizing it among negative ones, effectively optimizing mutual information [7]. Applying this framework to AU detection, however, faces two key challenges: severe class imbalance, where inactivated AUs dominate and bias predictions toward the negative class [8], and annotation noise, arising from subtle AU activations that complicate labeling and undermine model fitting [9].
To address these challenges, we propose a discriminative contrastive learning framework inspired by self-supervised learning for learning highly semantic image representations. This framework encodes differential information between AUs from each sample pair, providing an alternative to conventional pixel-level feature learning. The resulting AUNCE loss, conceptually similar to InfoNCE, is complemented by a negative sample re-weighting strategy [10] to mitigate class imbalance by assigning varying importance to individual samples. Additionally, a positive sample sampling strategy incorporates self-supervised signals via defined sampling ratios, ensuring that robust positive samples are derived through data augmentation.
Our research question is whether inter-AU differential encoding, combined with discriminative contrastive learning, can improve AU-specific representation learning and robustness under class imbalance and label noise. We hypothesize that such encoding enables more generalized and semantically consistent AU detection compared to conventional pixel-level feature learning approaches. Unlike prior purely self-supervised frameworks such as SimCLR [11], MoCo [12], KSRL [13], CLEF [14], et al., our approach integrates both supervised and self-supervised signals to balance semantic discrimination and representation robustness. Specifically, supervised AU labels (derived from manual annotations) enable the model to capture semantic category information instead of meaningless instance-level features—addressing a key limitation of purely self-supervised methods, which tend to overemphasize instance-specific differences. Meanwhile, self-supervised consistency signals (mined from data augmentations) enhance the model’s robustness against class imbalance and annotation noise. This joint design provides a principled means of more effectively leveraging limited yet noisy AU labels, thereby distinguishing our approach from existing contrastive-learning methods that lack such task-specific semantic guidance.
The primary contributions of our work are summarized as follows:
@@
    We introduce a positive sample sampling strategy to mitigate the effects of noisy and false AU labels.
@@ 2.1 Discriminative contrastive learning
In recent years, discriminative contrastive learning, a subset of deep self-supervised learning methods [11, 12], has witnessed substantial progress. This approach follows a learn-to-compare paradigm, distinguishing between similar (positive) and dissimilar (negative) data points in a feature space. For instance, Oord et al.[7] introduced the InfoNCE loss to optimize feature extraction by maximizing consistency in sequence data, establishing a foundation for contrastive methods. Chen et al.[11] proposed SimCLR, which uses a base encoder and projection head to maximize agreement through InfoNCE. Similarly, MoCo [12] leverages a momentum network to maintain a queue of negative samples for efficient learning. While variations in representation encoders and similarity measures exist, the core principle remains: training a representation encoder $f$ by contrasting similar pairs ($x,x^{+}$) against dissimilar pairs ($x,x^{-}$) to optimize the contrastive loss.
In addition to enabling the automatic extraction of fine-grained feature representations via the comparison of positive and negative pairs, deep contrastive learning has evolved to incorporate diverse contrastive objectives tailored for different supervision levels. In this context, Khosla et al. [15] extended the contrastive learning concept to supervised settings with a supervised contrastive learning loss, aiming to enhance feature expression using labeled data. Additionally, Robinson et al. [16] introduced a label-free strategy to sample hard negatives in contrastive learning, enabling tighter class clustering and improved downstream performance with minimal implementation overhead.
@@ 2.2 Contrastive learning for AU detection
Contrastive learning was firstly designed as a pretext task for unsupervised or self-supervised learning and was introduced to the domain of AU detection with limited annotations [17, 14]. For example, Niu et al. [17] introduced a semi-supervised co-training approach that utilized contrastive loss to enforce conditional independence between facial representations generated by two Deep Neural Networks (DNNs). Following this, Zhang et al. [14] introduced a weakly-supervised contrastive learning approach that effectively utilizes coarse-grained activity information.
@@
However, previous methods largely relied on weighted cross-entropy and triplet losses to encode pixel-level facial information, introducing redundant signals that can dilute AU-specific features. To mitigate this, Sun et al. [19] proposed EmoCo, an unsupervised contrastive framework using emotion-based image pairs; however, label inconsistencies and dependence on emotion annotations limit its applicability. Li et al. [9] employed a MoCo-based self-supervised approach for dynamic AU detection, yet its reliance on temporal cues restricts generalization to static images. Further, Li et al. [20] combined image- and feature-level contrastive learning to disentangle AU-relevant and domain-specific features, but performance may degrade under large domain shifts or limited AU diversity.
@@
Figure 2: Overview of the training pipeline of our discriminative contrastive learning framework. The training process consists of two stages: pretraining and linear evaluation. The pretraining stage precedes the linear evaluation stage. In the first stage, the feature encoder is pretrained by the AUNCE loss. In the second stage, we adopt a linear evaluation protocol, consistent with the practice established in contrastive learning frameworks such as SimCLR. Specifically, a single linear fully connected layer is trained atop the frozen encoder to assess the quality of the learned feature representations.
@@ 3.1 Objectives
To address these challenges, Section 3 provides detailed solutions. In Section 3.2, we propose a discriminative contrastive learning framework that directs the encoder to capture differential features between AUs from each sample pair, thus moving away from conventional pixel-level feature learning. To combat the class imbalance issue, Section 3.3.1 introduces a negative sample re-weighting strategy, drawing inspiration from, which prioritizes underrepresented AUs and facilitates the detection of challenging samples. Besides, Section 3.3.2 presents a positive sample sampling strategy that incorporates augmented supervision signals, which helps the model effectively learn from noisy and false data. Finally, Section 3.4 will introduce the encoder structure we used in our framework.
@@ 3.2 Discriminative contrastive learning framework
where $y_{i}$ denotes the ground truth value of occurrence (i.e., 0 or 1) of the $i$-th AU, and $\hat{y}_{i}$ is the corresponding predicted occurrence probability. This loss effectively guides the training process, encouraging the model to allocate higher probabilities to the correct class labels and lower probabilities to incorrect ones. $w_{i}$$\frac{(1/r_{i})n_{au}}{\sum_{i=1}^{n_{au}}(1/r_{i})}$ is the weight of the $i$-th AU determined by its occurrence frequency for balanced training. It permits reducing the impact of loss values associated with those AUs that are more commonly activated in the training set. $r_{i}$ represents the occurrence frequency of the $i$-th AU, and $n_{au}$ is the total number of AUs.
@@
To address this challenge, we introduce a discriminative contrastive learning framework for AU detection. As shown in Fig.2, the training process comprises two stages: pretraining and linear evaluation. The pretraining stage, preceding linear evaluation, integrates a feature encoder with two key components—a positive sample mining module and a negative sample re-weighting module. The former combines self-supervised and supervised signals to mitigate the impact of noisy labels caused by subtle AU variations, while the latter alleviates class imbalance and enhances hard negative sample mining. Further details are provided in Sections 3.3.1 and 3.3.2.
@@
The choice of ${L}_{aunce}$ over ${L}_{wce}$ in the pretraining phase is motivated by two key considerations. First, ${L}_{wce}$ relies on explicit label supervision and requires encoding pixel-level information to align with class labels, which imposes a heavier burden on the encoder. In contrast, ${L}_{aunce}$ operates on feature differences between sample pairs without requiring explicit labels, encouraging the encoder to capture discriminative cues between AUs more efficiently. Second, ${L}_{wce}$ constrains the absolute predicted probabilities, which may lead to overfitting and demand greater model capacity. Conversely, contrastive loss optimizes relative similarity between positive and negative pairs without enforcing exact values. This is because the minimal value of $\mathcal{L}_{aunce}$ is 0, which is achieved when $\text{sim}((f(x),f(x^{+}))-\text{sim}(f(x),f(x^{-}_{j}))\rightarrow+\infty$ for all negative sample index $j\in{1,2,\cdots,N}$.
In the linear evaluation phase, we follow a standard protocol commonly adopted in prior work [7, 11], where a linear classifier is trained atop the frozen encoder using weighted cross-entropy loss (WCE). Notably, this classifier is a single fully connected layer, which serves solely to evaluate representation quality, without influencing feature learning. Thus, the model’s performance primarily reflects the effectiveness of AUNCE in guiding representation learning.
3.3 Learning Strategies
3.3.1 Negative sample re-weighting strategy
@@
The key difference between our weighting strategy and that of HCL [16] is the omission of mean normalization in Eq. (3), allowing for more precise class imbalance mitigation in AU detection. We set $\beta\geq 1$ for majority class anchors to amplify minority class gradients, and $\beta\leq 1$ for minority class anchors to suppress majority class influence. By adjusting $\beta$, the model dynamically prioritizes underrepresented classes. Including mean normalization would mix $\beta$ with the normalization term, reducing interpretability and hindering gradient control, making our design critical for effective imbalance-aware learning.
3.3.2 Positive sample sampling strategy
To address the third challenge of AU labels tainted by noise and errors, we adopt a hybrid strategy that combines label noise learning with self-supervised techniques. As illustrated in Fig. 3, this method refines the supervision of positive samples by jointly leveraging self-supervised signal and supervised signals, thereby improving model robustness against annotation errors and inconsistencies.
@@ 3.3.2 Positive sample sampling strategy
Supervised Signal: Label smoothing is a widely used technique to mitigate label noise by introducing uncertainty through softened target distributions. However, its direct application in contrastive learning is non-trivial due to the absence of explicit labels in loss computation. To address this, we leverage the memory effect of neural networks—where models tend to learn clean samples before overfitting noisy ones. This phenomenon, observed across architectures, implies that positive samples with higher similarity to the anchor are more likely to be reliably learned.
Inspired by this observation, for a given sample $x$ and M positive samples ${x}_{i=1}^{M}$ with noisy labels belonging to the same class, we calculate the similarity scores between the representation of sample $x$ and the representations of all M positive samples. We then select the positive feature representation with the highest similarity to sample $x$:
@@
Self-supervised Signal: In addition to label noisy learning, we incorporate self-supervised learning techniques into the training process. By designing these auxiliary tasks carefully, we can encourage the model to learn more robust and invariant representations, which can help mitigate the effects of label noise and improve generalization performance. Specifically, our second method to obtain positive feature involves performing semantically invariant image augmentation on the anchor point $x$:
@@
Unlike data augmentation-based methods such as SimCLR [11], which rely on instance-level contrast using two augmented views of the same sample, our approach constructs positive examples as class prototypes—defined by the mean representation of multiple same-class samples. This design captures more stable and representative class-level semantics, enabling more robust feature comparisons. The resulting positive feature $f(x_{m}^{+})$, as defined in Eq. (10), serves as a virtual class centroid synthesized from M positive samples.
The three types of positive examples are constructed by combining label noise learning (for supervised signals) and data augmentation (for self-supervised signals), and are randomly sampled with probabilities $p_{1}$, $p_{2}$,and $p_{3}$ ($p_{1}+p_{2}+p_{3}=1$). This strategy enhances the quality of supervision during training and reduces the influence of noisy labels, thereby improving model robustness and generalization.
@@
Input: Mini batch data $\mathcal{R}$, feature representation function $f(\cdot)$, $p_{1}$, $p_{2}$, and $p_{3}$, $\beta$ .